# RegD: Hierarchical Embeddings via Distances over Geometric Regions

### v1

@@ Appendix E Ontologies
Ontologies use sets of statements (axioms) about concepts (unary predicates) and roles (binary predicates) for knowledge representation and reasoning. We focus on $\mathcal{EL}$-ontologies, which strike a good balance between expressivity and reasoning efficiency, making them widely applicable (Baader & Gil, 2024).
@@
An $\mathcal{EL}^{++}$-ontology is a finite set of TBox axioms of the form
@@
[…]
@@
if $C\sqsubseteq D$ is satisfied by all models of $\mathcal{O}$. Ontologies can infer much more complex patterns than the transitive closure.
@@
[…]
Appendix F Experiments Setting
Hyperparameters
In our experiments, we explored various hyperparameter configurations, focusing primarily on performance in low-dimensional spaces. Following the setup in (Yu et al., 2024), we evaluated our models in dimensions $5$ and $10$. The weighting coefficient $\lambda$ in Eq. 7 was selected from $\{0,0.5,1\}$, while the learning rate was tuned over $\{0.01,0.005,0.001,0.0005\}$. For the small dataset of mammals, which contains approximately 1000 nodes, we choose batch sizes from $\{32,64,128\}$. For the remaining datasets, we opt for batch sizes from $\{1024,2048,8192\}$. For the depth distance, we tested $p\in\{1,2\}$ as specified in Definition 2. The boundary distance margins in Eq. 4 were set to $\gamma_{1}=0.001$ and $\gamma_{2}=0$ for directed acyclic graph (DAG) embedding tasks. In ontology embedding experiments, both margins were set to zero.
For DAG experiments, as in (Yu et al., 2024), we trained for 400 epochs on the mammal and noun datasets and 500 epochs on the MCG and Hearst datasets. For ontology embedding experiments, we trained for 5000 epochs, following the protocol in previous work (Jackermeier et al., 2024).
Ablation Studies
The results of the ablation studies are summarized in Table 5. Note that, when $\lambda=0$, the depth distance component is absent from the energy function, and therefore, the choice of $p$, as a hyperparameter of depth distance, is not applicable.
From the results, it is evident that our boundary distance generally outperforms the specific boundary distances for boxes and balls (i.e., $d_{\textit{vol}}$ and $d_{\text{bd}}^{\text{cone}}$) in most cases. Furthermore, when using boxes as the embedding object, setting $p=1$ yields better results than $p=2$. However, for balls, $p=2$ achieves superior performance.
For the mammal and noun datasets, using balls as the embedding object and incorporating the depth distance (i.e., $\lambda>0$) in the energy function results in better performance. Conversely, for the MCG and Hearst datasets, the best performance is typically achieved when using boxes as the embedding region and setting $\lambda=0$, effectively excluding the hyperbolic components. This observation aligns with our analysis in the main text.
It is worth noting, as shown in the seventh row of Table 5, that while the results using the depth distance are not always the best, they are very close to the best performance. This suggests that our depth distance configuration has the potential to achieve even better results with further refinement.
Options Mammal Noun MCG Hearst
region $\lambda$ p boundary distance d=2 d=5 d=5 d=10 d=5 d=10 d=5 d=10
Box 0 - $d_{\text{bd}}$ 39.4 42.8 34.6 34.6 49.9 56.6 41.6 44.7
$d_{\textit{vol}}$ 31.7 38.1 30.8 31.9 48.6 58.5 42.8 49.6
$>0$ 1 $d_{\text{bd}}$ 64.9 74.4 53.8 51.3 43.7 45.0 35.6 38.8
$d_{\textit{vol}}$ 45.6 54.3 30.4 33.4 30.9 39.9 32.6 38.8
2 $d_{\text{bd}}$ 56.1 62.2 50.3 47.4 50.7 57.4 40.6 44.6
$d_{\textit{vol}}$ 39.7 47.9 30.6 31.6 48.5 56.7 37.8 43.8
Ball 0 - $d_{\text{bd}}$ 42.8 55.5 39.6 42.5 43.9 46.7 37.7 37.1
$d_{\text{bd}}^{\text{cone}}$ 41.2 49.7 36.4 37.7 40.9 43.7 36.0 36.7
$>0$ 1 $d_{\text{bd}}$ 58.2 69.2 51.7 54.4 40.9 45.9 33.6 35.8
$d_{\text{bd}}^{\text{cone}}$ 57.9 65.3 52.3 54.7 36.8 41.7 31.6 34.6
2 $d_{\text{bd}}$ 62.5 71.8 58.4 59.1 41.7 45.8 34.8 37.7
$d_{\text{bd}}^{\text{cone}}$ 62.7 67.7 54.3 55.1 40.7 45.7 34.6 36.7
F.1 Other Results
Including non-basic edges in training
Table 6: F1 score (%) on mammal sub-graph with best numbers bolded. Results with * are coming from Yu et al. (2024).
Non-basic-edge Percentage Dimension = 2 Dimension = 5
0% 10% 25% 50% 90% 0% 10% 25% 50% 90%
GBC-box 23.4 25.0 23.7 43.1 48.2 35.8 60.1 66.8 83.8 97.6
VBC-box 20.1 26.1 31.0 33.3 34.7 30.9 43.1 58.6 74.9 69.3
$\tau$Box 29.0 33.0 41.2 49.6 53.5 33.5 37.5 45.0 58.9 64.0
Entailment Cone 54.4 61.0 71.0 66.5 73.1 56.3 81.0 84.1 83.6 82.9
ShadowCone (Umbral-half-space) 57.7 73.7 77.4 80.3 79.0 69.4 81.1 83.7 88.5 91.8
Ours (box) 64.9 74.4 75.8 78.3 82.8 71.6 77.8 83.2 87.6 87.0
Ours (ball) 62.7 68.3 77.2 84.1 88.6 71.8 74.5 84.4 88.4 90.2
Using learnable function $h$ in Equations 2 and 3
We investigated alternative formulations of the function $h$ beyond using a constant value. Specifically, we evaluated two approaches:
  1. 1.
    $h(\text{{reg}})=1/f_{\text{MLP}}([\text{para}(\text{{reg}}_{1})])$
  2. 2.
    $h(\text{{reg}})=f_{\text{MLP}}([\text{para}(\text{{reg}}_{1})])$
where $f_{\text{MLP}}$ denotes a two-layer neural network defined as:
$f_{\text{MLP}}(x)=\text{SoftPlus}\left(W^{(2)}\,\text{ReLU}(W^{(1)}x+b^{(1)})+b^{(2)}\right)$
The SoftPlus activation function ensures the positivity of $h(\text{{reg}})$. In the implementation, we only use the center of balls and boxes as the input of $h$.
Our experimental results demonstrated that a constant value of $h(\text{{reg}})=1$ generally achieved optimal performance across most scenarios. Neural network-based formulations of $h(\text{{reg}})$ showed advantages only in specific cases, such as with the heart dataset. For instance, when $d=5$, $p=2$, and $\lambda>0$, using $h(\text{{reg}})=1/f_{\text{MLP}}$ improved the F1-score from 34.8 to 37.7 (see Table 5, intersection of the second-to-last row and second-to-last column). These findings suggest that while adaptive formulations of $h(\text{{reg}})$ can offer marginal improvements in specific scenarios, they require careful consideration in both design and training. For hierarchy-specific tasks, the simpler constant formulation of $h(\text{{reg}})$ typically provides sufficient and stable performance.
F.2 Ontology embedding methods
@@
    Note that the original ELEM method includes a regularization term that enforces the norms $\|\mathbf{c}_{u}\|$ and $\|\mathbf{c}_{v}\|$ to be close to 1. However, for the DAG case, we omit this regularization primarily applies to scenarios involving relation embeddings, such as axioms of the form $A\sqsubseteq\exists r.B$ or $\exists r.B\sqsubseteq A$, as in the KGE methods TransE (Bordes et al., 2013).
Integration of ELBE and ELEM with RegD and $\tau$Box
@@
To integrate RegD and $\tau$Box into ELBE or ELEM, it suffices to replace their energy function for axioms $C\sqsubseteq D$ with the corresponding energy function from RegD or $\tau$Box. However, ELBE and ELEM handle negative samples differently, using a loss function of the form:
@@
When integrating with RegD, this loss function should also be replaced with the one defined in Section 5.

### v2

@@ Appendix E Ontologies
Ontologies use sets of statements (axioms) about concepts (unary predicates) and roles (binary predicates) for knowledge representation and reasoning. We focus on $\mathcal{EL}$-ontologies, which strike a good balance between expressivity and reasoning efficiency, making them widely applicable Baader & Gil (2024).
@@
An $\mathcal{EL}$-ontology is a finite set of TBox axioms of the form
@@
$\text{Father}\sqsubseteq\text{Male}\sqcap\text{Parent},\quad\text{Child}\sqsubseteq\exists\text{hasParent.Father}.$
@@
if $C\sqsubseteq D$ is satisfied by all models of $\mathcal{O}$. Ontologies can infer much more complex patterns than the transitive closure, as shown by the following example.
@@
In an ontology, suppose that in a group $X$, all members are both a man and a parent. That is, $X\sqsubseteq\textit{man}$ and $X\sqsubseteq\textit{parent}$ in the ontology. Moreover, man and parent together imply father (formally, $\textit{man}\sqcap\textit{parent}\sqsubseteq\textit{father}$ in the ontology). From this, we can infer that the group $X$ is a father, which is not derivable from the transitive closure.
E.1 Ontology embedding methods
@@
    Note that the original ELEM method includes a regularization term that enforces the norms $\|\mathbf{c}_{u}\|$ and $\|\mathbf{c}_{v}\|$ to be close to 1. However, for the DAG case, we omit this regularization primarily applies to scenarios involving relation embeddings, such as axioms of the form $A\sqsubseteq\exists r.B$ or $\exists r.B\sqsubseteq A$, as in the KGE methods TransE Bordes et al. (2013).
[…]
@@
[…]
@@
When integrating with RegD, this loss function should also be replaced with the one defined in Section 3.3.
Result of prediction task by ranking-based metrics
The evaluation results for the prediction task on ontologies are presented in Table 9. RegD shows consistent improvements for both ELEM and ELBE across most cases. The gains are particularly notable for the GO and ANATOMY ontologies, where the H@100 score increases from 0 to 0.26. In contrast, performance on the GALEN ontology remains low for both variants (with or without RegD), likely due to the inherent complexity of GALEN. This suggests that more advanced methods may be required to effectively model such complex structures.
Appendix F The Use of LLMs
LLMs have been employed solely to assist with polishing the writing. All content was originally written by the authors and subsequently refined by LLMs to improve readability, clarity, and to correct grammatical errors or typographical mistakes.
Table 9: Prediction task
Dataset Method H@1 H@10 H@100 Medain MRR MR AUC
GALEN ELBE 0 0 0 7831 0 9044 0.61
(d=5) + RegD 0 0 0.02 9506 0 10056 0.57
ELEM 0 0.01 0.04 8843 0 9361 0.60
+ RegD 0 0 0.02 10122 0 10531 0.55
GALEN ELBE 0 0 0 8812 0 $95$ 0.59
(d=10) + RegD 0 0.01 0.02 8966 0 9729 0.58
ELEM 0 0 0.03 8856 0 9534 0.59
+ RegD 0 0.01 0.03 9996 0 10441 0.55
GO ELBE 0 0 0 14058 0 17231 0.62
(d=5) + RegD 0 0.10 0.19 8466 0.02 14376 0.69
ELEM 0 0 0.01 19666 0 19813 0.57
+ RegD 0 0.01 0.03 15440 0 17983 0.61
GO ELBE 0 0 0 18926 0 19656 0.57
(d=10) + RegD 0 0.17 0.26 7785 0.04 14244 0.69
ELEM 0 0 0.01 19668 0 20543 0.55
+ RegD 0 0.02 0.06 17194 0.01 19019 0.59
ANATOMY ELBE 0 0 0.01 19556 0 29528 0.72
(d=5) + RegD 0 0.02 0.05 14317 0.01 25582 0.76
ELEM 0 0.01 0.04 23437 0 32788 0.69
+ RegD 0 0.04 0.14 8616 0.01 27272 0.74
ANATOMY ELBE 0 0 0.03 14955 0 25844 0.76
(d=10) + RegD 0 0.05 0.13 10019 0.02 22215 0.79
ELEM 0 0.01 0.06 14573 0 27010 0.75
+ RegD 0 0.07 0.22 8066 0.03 27632 0.74